If P is a point of intersection of the diagonals of the parallelogram ABCD If O is any point then OA+OB+OC+OD=

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

we know digonal of parallelogram intersect each other

so P is the mid pont of A and C and also p is the mid point of B and D

OP = (OA + OC )/2   or 2OP = OA + OC

and

OP = (OB + OD  )/2  or 2OP = OB + OD

add

4OP =OA+OB+OC+OD

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin